OpenStudy (anonymous):

Which value of y makes the equation true? y – 6.3 = –2.2 y =

OpenStudy (anonymous):

If you isolate y, you will also get the value.

OpenStudy (anonymous):

To have y alone, add +6.3 to both sides.

OpenStudy (anonymous):

y-6.3+6.3=-2.2+6.3 y=4.1

OpenStudy (anonymous):

Let's check if the value gives a true statement: 4.1-6.3=-2.2 -2.2=-2.2 It does.

OpenStudy (anonymous):

Do you want me to give you one to try on your own?